# Early stopping for conjugate gradients in statistical inverse problems

### v1

@@ 7.2 Early stopping and its reconstruction error
We analyse the reconstruction error of the early stopping rule $\tau$ from Definition 3.4. Unfortunately, there is no reconstruction error analogue of the oracle inequality in Theorem 6.9. Instead, we need to directly bound $\lvert r_{\tau}^{\prime}(0)\rvert$ and the errors at $\tau$ in high probability to apply Corollary 7.3. For this purpose Lemmas 6.5 and 6.7 are essential, requiring $\lvert r_{\tau}^{\prime}(0)\rvert<\lambda_{D}^{-2}$. The case $\lvert r_{\tau}^{\prime}(0)\rvert\geqslant\lambda_{D}^{-2}$ leads to high stochastic error and is dealt with separately, assuming the same lower bound on the dimension $D$ as for the lower bound on the reconstruction error in Remark 7.7. The proofs for this subsection are delegated to Section 9.5.
@@
Assume (GN), (PSD($p$,$c_{A}$,$C_{A}$)) and put $M_{\tau,\lambda}\coloneqq A_{\tau,\lambda}\vee S_{\tau,\lambda}$. Under the source condition $\lVert f^{\dagger}\rVert_{\mu}\leqslant R$, the following holds:
@@
Assume (PSD($p$,$c_{A}$,$C_{A}$)) with $p>1/2$, (GN) and $\lvert\kappa-\delta^{2}D\rvert\leqslant C_{\kappa}\delta^{2}\sqrt{D}$ for some $C_{\kappa}\geqslant 0$. Under the source condition $\lVert f^{\dagger}\rVert_{\mu}\leqslant R$, we have for all $z\geqslant 1$ with probability at least $1-6e^{-z}$
@@
provided $(R^{2}\delta^{-2})^{1/(4\mu p+2p+1)}\lesssim D$. Here $C_{\mu,p,c_{A},C_{A},C_{\kappa}},\widetilde{C}_{\mu,p,c_{A},C_{A},C_{\kappa}}$ are positive constants depending only on $\mu$, $p$, $c_{A}$, $C_{A}$ and $C_{\kappa}$. This gives the minimax rate for all regularities $\mu>0$ such that $\sqrt{D}\lesssim(R^{2}\delta^{-2})^{1/(4\mu p+2p+1)}\lesssim_{\mu,p,c_{A}}D$.
@@
Under (DN) and the source condition $\lVert f^{\dagger}\rVert_{\mu}\leqslant R$, the discrepancy principle $\tau$ with $\kappa\coloneqq c^{2}\delta^{2}$, $c>1$, satisfies
@@ 8 Numerical examples
In both settings, we calculate the conjugate gradient estimator $\widehat{f}_{\tau}$, where the threshold for the early stopping rule is chosen as $\kappa\coloneqq\delta^{2}D$ for the first three examples and $\kappa\coloneqq\delta^{2}D+\delta^{2}\sqrt{D}$ for the gravity problem. At the latest, the algorithm is terminated at step $T\in{\mathbb{N}}$ if $T$ reaches the maximal iteration index $D$ or if the squared norm $\lVert A^{\top}(Y-A\widehat{f}_{T})\rVert^{2}$ of the transformed residual vector is smaller than $10^{-8}$ (emergency stop). Otherwise, it would break down with division by zero [8, p. 179]. Note that for ill-conditioned matrices, $R_{t}^{2}$ does not have to be close to zero for $\lVert A^{\top}(Y-A\widehat{f}_{t})\rVert^{2}$ to be small.
@@
for the prediction and reconstruction error, respectively. Ideally, the values should be close to one.
@@
Remark 7.9 explains that we cannot expect $\lvert S_{\tau,\lambda}-A_{\tau,\lambda}\rvert$ to be small for severely ill-posed problems so that we loose control of $\lvert r_{\tau}^{\prime}(0)\rvert$. Therefore, we recommend to stop slightly earlier and we choose here $\kappa\coloneqq\delta^{2}D+\delta^{2}\sqrt{D}$, still satisfying our condition on $\kappa$ in Corollary 6.10 with $C_{\kappa}=1$, leading to emergency stops in $21$ percent of the runs. If we ignore these runs, the minimal relative efficiency for the reconstruction error is about $0.2$ such that in the worst case the reconstruction error is five times the optimal error. The interesting conclusion is that the early stopping criterion provides reasonable results even in a setting not covered by the theory.
@@
For all three signals and both error criteria, the early stopped estimators closely match the optimal ones. For the rough signal, the asymptotic regime applies only for small $\delta$, that is large $D$. This is explained by the fact that the signals still change with $D$, having more nonzero coefficients and thus inducing larger approximation errors. Again, the similarity is even higher for the reconstruction error. The slopes of the minimax rates are well matched, even exhibiting a small asymptotic constant.
@@ 9.2 Results for classical conjugate gradients
Recall Definition 3.1 of the noninterpolated estimators. For $k<d$, the minimising polynomial $r_{k}$ is given by the orthogonal projection of the origin on the affine subspace $\operatorname{Pol}_{k,1}$ of $\operatorname{Pol}_{k}$ and is therefore unique. In the case $k=d$, the zeros of $r_{d}$ are given by the distinct squared nonzero singular values of $A$ as shown in 9.5(a). Since $r_{d}(0)=1$, $r_{d}$ is also uniquely determined.
@@
First, consider the case $k<d$. By the definition of $r_{k}$, the polynomial $r_{k}-1$ is the orthogonal projection of $-1$ with respect to $\lVert\cdot\rVert_{Y}$ onto the linear subspace $\operatorname{Pol}_{k,0}$. Therefore, $r_{k}=r_{k}-1-(-1)$ is orthogonal to $\operatorname{Pol}_{k,0}$ with respect to $\langle{\cdot},{\cdot}\rangle_{Y}$. For $k=d$, the claim follows from 9.5(a). ∎
@@
where $\lVert\mathsf{x^{1/2}r_{l}}\rVert_{Y}>0$ for all $l=0,\dots,d-1$.
@@
Without loss of generality, assume $l<k$. Since $x\mapsto xr_{l}(x)\in\operatorname{Pol}_{l+1,0}\subset\operatorname{Pol}_{k,0}$, Lemma 9.3 yields $\langle{\mathsf{r_{k}}},{\mathsf{xr_{l}}}\rangle_{Y}=0$. For $l<d$, there is $i^{*}\in\{1,\dots,d\}$ such that $r_{l}(\tilde{\lambda}_{i^{*}}^{2})\neq 0$, since a polynomial of degree $l$ can have at most $l$ zeros. Hence, by (Y), $\lVert\mathsf{x^{1/2}r_{l}}\rVert_{Y}^{2}\geqslant\tilde{\lambda}_{i^{*}}^{2}r_{l}(\tilde{\lambda}_{i^{*}}^{2})^{2}\sum_{j:\,\lambda_{j}=\tilde{\lambda}_{i^{*}}}Y_{j}^{2}>0$, which concludes the proof. ∎
@@
    $r_{k}$ has $k$ real simple zeros $0<x_{1,k}<\dots<x_{k,k}<\lVert AA^{\top}\rVert_{\mathrm{spec}}$, $k=1,\dots,d-1$. For $k=d$, the zeros are the $d$ distinct squared nonzero singular values $0<\tilde{\lambda}_{d}^{2}<\dots<\tilde{\lambda}_{1}^{2}=\lVert AA^{\top}\rVert_{\mathrm{spec}}$.
@@
    Corollary 9.4 shows that $(r_{k})_{k=0,\dots,d}$ is orthogonal in $L^{2}([0,\lVert AA^{\top}\rVert_{\mathrm{spec}}],\hat{\mu})$ with $\hat{\mu}\coloneqq\sum_{i=1}^{D}\lambda_{i}^{2}Y_{i}^{2}\delta_{\lambda_{i}^{2}}$ and $\lVert r_{k}\rVert_{L^{2}(\hat{\mu})}>0$ for $k=0,\dots,d-1$. The first part for $k<d$ is a well-known fact for orthogonal polynomials, see Szegő [24, Theorem 3.3.1]. Since $A$ has $d$ distinct nonzero singular values, we have $\lVert\mathsf{r_{d}}\rVert_{Y}^{2}=\min_{p\in\operatorname{Pol}_{d,1}}\lVert\mathsf{p}\rVert_{Y}^{2}=0$. By Assumption (Y), this implies $r_{d}(\tilde{\lambda}_{i}^{2})=0$, $i=1,\dots,d$, and we conclude the second part.
@@
    $x_{k,k}<x_{d,d}=\tilde{\lambda}_{1}^{2}=\lambda_{1}^{2}.$
    In particular, the claim holds for $k=1$. For $k=2,\dots,d-1$, fix $j\in\{1,\dots,d-1\}$ and assume that there exists an $l\in\{1,\dots,k-1\}$ such that $x_{l,k},x_{l+1,k}\in(\tilde{\lambda}_{j+1}^{2},\tilde{\lambda}_{j}^{2}]$. Let
    $f(x)\coloneqq\Big{(}\prod_{\begin{subarray}{c}i=1,\dots,k,\\ i\neq l,l+1\end{subarray}}(x-x_{i,k})^{2}\Big{)}(x-x_{l,k})(x-x_{l+1,k}),\quad x\in{\mathbb{R}}.$
    For $x\in\{\tilde{\lambda}_{1}^{2},\dots,\tilde{\lambda}_{d}^{2}\}$, there are two cases: if $x\leqslant\tilde{\lambda}_{j+1}^{2}$, then $x-x_{l+1,k}<x-x_{l,k}<0$; otherwise, if $x\geqslant\tilde{\lambda}_{j}^{2}$, then $0\leqslant x-x_{l+1,k}<x-x_{l,k}$. In both cases, we have
    $(x-x_{l,k})(x-x_{l+1,k})\geqslant 0.$
    Thus, $f(x)\geqslant 0$ for all $x\in\{\tilde{\lambda}_{1}^{2},\dots,\tilde{\lambda}_{d}^{2}\}$. In particular, $x_{i,k}<x_{k,k}<\tilde{\lambda}_{1}^{2}$, $i=1,\dots,k-1$, implies $f(\tilde{\lambda}_{1}^{2})>0$. Hence, under (Y),
    $\sum_{i=1}^{D}\lambda_{i}^{2}f(\lambda_{i}^{2})Y_{i}^{2}\geqslant\tilde{\lambda}_{1}^{2}f(\tilde{\lambda}_{1}^{2})\sum_{j:\,\lambda_{j}=\tilde{\lambda}_{1}}Y_{j}^{2}>0.$
    On the other hand, we can write
    $f(x)=c^{-1}r_{k}(x)\cdot s_{k-2}(x)\quad\text{with}\quad s_{k-2}(x)\coloneqq\prod_{\begin{subarray}{c}i=1,\dots,k,\\ i\neq l,l+1\end{subarray}}(x-x_{i,k})\in\operatorname{Pol}_{k-2},$
    where $c\coloneqq\prod_{i=1}^{k}-x_{i,k}^{-1}$. We deduce
    $\sum_{i=1}^{D}\lambda_{i}^{2}f(\lambda_{i}^{2})Y_{i}^{2}=\sum_{i=1}^{D}r_{k}(\lambda_{i}^{2})q_{k-1}(\lambda_{i}^{2})Y_{i}^{2}=\langle{\mathsf{r_{k}}},{\mathsf{q_{k-1}}}\rangle_{Y},$
    where $q_{k-1}(x)\coloneqq c^{-1}xs_{k-2}(x)\in\operatorname{Pol}_{k-1,0}$. By Lemma 9.3, we have $\langle{\mathsf{r_{k}}},{\mathsf{q_{k-1}}}\rangle_{Y}=0$, which yields a contradiction. We conclude that every interval $(\tilde{\lambda}_{j+1}^{2},\tilde{\lambda}_{j}^{2}]$, $j=1,\dots,d-1$, contains at most one zero of $r_{k}$. Since $x_{k,k}<\tilde{\lambda}_{1}^{2}$, this implies $x_{k+1-i,k}\leqslant\tilde{\lambda}_{i}^{2}\leqslant\lambda_{i}^{2}$ for $k=1,\dots,d$, $i=1,\dots,k$. ∎
The following result is originally due to Nemirovskii [20, (3.14)], who introduced the function $\varphi_{k}^{2}$, named $r_{k}$ therein. The proof of the first inequality can also be found in Hanke [10, (3.8)] and Engl, Hanke and Neubauer [8, (7.7)]. For the sake of completeness, we state it in our notation below. The second inequality is new, but follows easily from 9.5(b) and the definition of $\varphi_{k}$.
@@ 9.3 Proofs for Section 4
    The definitions of $r_{t_{1}}$ and $r_{t_{2}}$ and (9.2) yield
@@
    since $r_{k+1}(x)<r_{k}(x)$ for all $x\in(0,x_{1,k+1}]$ by 9.5(e), $r_{k}(x)\geqslant 0$ for $x\in[0,x_{1,k}]$ and $r_{k+1}(x)<0$ for $x\in(x_{1,k+1},x_{1,k}]$. This implies $x_{1,t_{2}}<x_{1,t_{1}}$. Let $x>x_{k+1,k+1}$. For $k$ even we obtain $r_{k}(x)>0$ and $r_{k+1}(x)<0$ such that $r_{t_{1}}(x)>r_{t_{2}}(x)$. Since $r_{t_{j}}(x_{k+1,k+1})>0$, $j=1,2$, this yields $x_{k+1,t_{2}}<x_{k+1,t_{1}}$. In the same way, for $k$ odd we deduce $r_{t_{1}}(x)<r_{t_{2}}(x)$ and $r_{t_{j}}(x_{k+1,k+1})<0$, $j=1,2$, implying $x_{k+1,t_{2}}<x_{k+1,t_{1}}$.
@@
and note the order $0\leqslant\tilde{x}_{1,t}<x_{1,k+1}<x_{1,t}$ due to $x_{i,t}>x_{i,k+1}$ by 4.5(b). With some $q_{k},\tilde{q}_{k},\check{q}_{k}\in\operatorname{Pol}_{k,0}$, $\tilde{p}_{k-1}\in\operatorname{Pol}_{k-1,1}$, we have
@@
$=\frac{\tilde{x}_{1,t}}{x_{1,t}}+\check{q}_{k}(x)-\frac{x}{x_{1,t}}\bigg{(}x^{k}\prod_{i=2}^{k+1}\Big{(}-\frac{1}{x_{i,t}}\Big{)}+\tilde{p}_{k-1}(x)\bigg{)}$
@@
$\leqslant\Big{(}\frac{2\mu+1}{e\lvert r_{t}^{\prime}(0)\rvert}\Big{)}^{2\mu+1}\lVert g\rVert_{\mu}^{2}$

### v2

@@ 7.2 Early stopping and its reconstruction error
We analyse the reconstruction error of the early stopping rule $\tau$ from Definition 3.3. Unfortunately, there is no reconstruction error analogue of the oracle inequality in Theorem 6.8. Instead, we need to directly bound $\lvert r_{\tau}^{\prime}(0)\rvert$ and the errors at $\tau$ in high probability to apply Corollary 7.3. For this purpose Lemmas 6.4 and 6.6 are essential, requiring $\lvert r_{\tau}^{\prime}(0)\rvert<\lambda_{D}^{-2}$. The case $\lvert r_{\tau}^{\prime}(0)\rvert\geqslant\lambda_{D}^{-2}$ leads to high stochastic error and is dealt with separately, assuming the same lower bound on the dimension $D$ as for the lower bound on the reconstruction error in Remark 7.7. The proofs for this subsection are delegated to Section 9.5.
@@
Assume (GN), (PSD($p$,$c_{A}$,$C_{A}$)) and put $M_{\tau,\lambda}\coloneqq A_{\tau,\lambda}\vee S_{\tau,\lambda}$. Under the source condition $\lVert f\rVert_{\mu}\leqslant R$, the following holds:
@@
Assume (PSD($p$,$c_{A}$,$C_{A}$)) with $p>1/2$, (GN) and $\lvert\kappa-\delta^{2}D\rvert\leqslant C_{\kappa}\delta^{2}\sqrt{D}$ for some $C_{\kappa}\geqslant 0$. Under the source condition $\lVert f\rVert_{\mu}\leqslant R$, we have for all $z\geqslant 1$ with probability at least $1-6e^{-z}$
@@
provided $(R^{2}\delta^{-2})^{1/(4\mu p+2p+1)}\lesssim D$, where $\mathcal{R}^{*,\mathrm{rec}}_{\mu,p,R}(\delta)$ is given in (3.5). Here $C_{\mu,p,c_{A},C_{A},C_{\kappa}}$, $\widetilde{C}_{\mu,p,c_{A},C_{A},C_{\kappa}}$ are positive constants depending only on $\mu$, $p$, $c_{A}$, $C_{A}$ and $C_{\kappa}$. This gives the minimax rate for all regularities $\mu>0$ such that $\sqrt{D}\lesssim(R^{2}\delta^{-2})^{1/(4\mu p+2p+1)}\lesssim_{\mu,p,c_{A}}D$.
@@
Under (DN) and the source condition $\lVert f\rVert_{\mu}\leqslant R$, the discrepancy principle $\tau$ with $\kappa\coloneqq c^{2}\delta^{2}$, $c>1$, satisfies
@@ 8 Numerical examples
In both settings, we calculate the conjugate gradient estimator $\widehat{f}_{\tau}$, where the threshold for the early stopping rule is chosen as $\kappa\coloneqq\delta^{2}D$ for the first three examples and $\kappa\coloneqq\delta^{2}D+\delta^{2}\sqrt{D}$ for the gravity problem. At the latest, the algorithm is terminated at step $T\in{\mathbb{N}}$ if $T$ reaches the maximal iteration index $D$ or if the squared norm $\lVert A^{\top}(Y-A\widehat{f}_{T})\rVert^{2}$ of the transformed residual vector is smaller than $10^{-8}$ (emergency stop). Otherwise, it would break down with division by zero [8, p. 179]. Note that for ill-conditioned matrices, $R_{t}^{2}$ is not necessarily close to zero for $\lVert A^{\top}(Y-A\widehat{f}_{t})\rVert^{2}$ to be small.
@@
for the prediction and reconstruction error, respectively. The values lie in $[0,1]$ and ideally should be close to one.
@@
Remark 7.9 explains that we cannot expect $\lvert S_{\tau,\lambda}-A_{\tau,\lambda}\rvert$ to be small for severely ill-posed problems so that we loose control of $\lvert r_{\tau}^{\prime}(0)\rvert$. Therefore, we recommend to stop slightly earlier and we choose here $\kappa\coloneqq\delta^{2}D+\delta^{2}\sqrt{D}$, still satisfying our condition on $\kappa$ in Corollary 6.9 with $C_{\kappa}=1$, leading to emergency stops in $21$ percent of the runs. If we ignore these runs, the minimal relative efficiency for the reconstruction error is about $0.13$ such that in the worst case the reconstruction error is eight times the optimal error. The interesting conclusion is that the early stopping criterion provides reasonable results even in a setting not covered by the theory.
@@
For all three signals and both error criteria, the early stopped estimators closely match the optimal ones. For the rough signal, the asymptotic regime applies only for small $\delta$, that is large $D$. This is explained by the fact that the signals still change with $D$, having more nonzero coefficients and thus inducing larger approximation errors. Again, the similarity is even higher for the reconstruction error. Note that the choice $\mu=1/4$ corresponds to the rough signal, for which the slopes of the minimax rates are well-matched. For the smoother signals, the convergence is faster, as expected for an adaptive method.
@@ 9.2 Results for classical conjugate gradients
Recall Definition 3.1 of the noninterpolated estimators.
Lemma 9.3.
The residual polynomial $r_{k}$ is well-defined for $k=0,\dots,d$. Moreover, $r_{d}$ is uniquely determined by its zeros $0<\tilde{\lambda}_{d}^{2}<\dots<\tilde{\lambda}_{1}^{2}=\lVert AA^{\top}\rVert_{\mathrm{spec}}$.
Proof.
For $k<d$, the orthogonal projection with respect to $\langle{\cdot},{\cdot}\rangle_{Y}$ of the origin on the affine subspace $\operatorname{Pol}_{k,1}$ of $\operatorname{Pol}_{k}$ uniquely solves the minimisation problem $\min_{p_{k}\in\operatorname{Pol}_{k,1}}\lVert\mathsf{p_{k}}\rVert_{Y}^{2}$. Therefore, $r_{k}$ is well-defined for $k<d$. In the case $k=d$, note that $\min_{p_{d}\in\operatorname{Pol}_{d,1}}\lVert\mathsf{p_{d}}\rVert_{Y}^{2}=0$, since $A$ has $d$ distinct nonzero singular values. By Assumption (Y), the solution is uniquely defined, namely as the polynomial in $\operatorname{Pol}_{d,1}$, which has the squared singular values $\tilde{\lambda}_{i}^{2}$, $i=1,\dots,d$, as zeros. This proves the well-definedness and claimed characterisation of $r_{d}$. ∎
@@
First, consider the case $k<d$. By the definition of $r_{k}$, the polynomial $r_{k}-1$ is the orthogonal projection with respect to $\langle{\cdot},{\cdot}\rangle_{Y}$ of $-1$ onto the linear subspace $\operatorname{Pol}_{k,0}$. Therefore, $r_{k}=r_{k}-1-(-1)$ is orthogonal to $\operatorname{Pol}_{k,0}$ with respect to $\langle{\cdot},{\cdot}\rangle_{Y}$. For $k=d$, the claim follows from Lemma 9.3. ∎
@@
where $\lVert\mathsf{x^{1/2}r_{l}}\rVert_{Y}>0$ for all $l=0,\dots,d-1$. This implies that $r_{k}$ has degree $k$ for $k=0,\dots,d$.
@@
Without loss of generality, assume $l<k$. Since $x\mapsto xr_{l}(x)\in\operatorname{Pol}_{l+1,0}\subset\operatorname{Pol}_{k,0}$, Lemma 9.4 yields $\langle{\mathsf{r_{k}}},{\mathsf{xr_{l}}}\rangle_{Y}=0$. For $l<d$, there is $i^{*}\in\{1,\dots,d\}$ such that $r_{l}(\tilde{\lambda}_{i^{*}}^{2})\neq 0$, since a polynomial of degree at most $l$ can have at most $l$ zeros. Hence, by Assumption (Y), $\lVert\mathsf{x^{1/2}r_{l}}\rVert_{Y}^{2}\geqslant\tilde{\lambda}_{i^{*}}^{2}r_{l}(\tilde{\lambda}_{i^{*}}^{2})^{2}\sum_{j:\,\lambda_{j}=\tilde{\lambda}_{i^{*}}}Y_{j}^{2}>0$. For the second part, assume $r_{k}$ has degree $l<k$. Then, by Lemma 9.3, we have $r_{k}=r_{l}$ such that $\langle{\mathsf{r_{k}}},{\mathsf{xr_{l}}}\rangle_{Y}=\lVert\mathsf{x^{1/2}r_{l}}\rVert_{Y}^{2}>0$. On the other hand, we have shown $\langle{\mathsf{r_{k}}},{\mathsf{xr_{l}}}\rangle_{Y}=0$, which yields a contradiction. This concludes the proof. ∎
@@
    $r_{k}$ has $k$ real simple zeros $0<x_{1,k}<\dots<x_{k,k}<\lVert AA^{\top}\rVert_{\mathrm{spec}}$, $k=1,\dots,d-1$. For $k=d$, the zeros are the $d$ distinct squared nonzero singular values $0<x_{1,d}=\tilde{\lambda}_{d}^{2}<\dots<x_{d,d}=\tilde{\lambda}_{1}^{2}=\lVert AA^{\top}\rVert_{\mathrm{spec}}$.
@@
    Corollary 9.5 shows that $(r_{k})_{k=0,\dots,d}$ is orthogonal in $L^{2}([0,\lVert AA^{\top}\rVert_{\mathrm{spec}}],\hat{\mu})$ with $\hat{\mu}\coloneqq\sum_{i=1}^{D}\lambda_{i}^{2}Y_{i}^{2}\delta_{\lambda_{i}^{2}}$ and $\lVert r_{k}\rVert_{L^{2}(\hat{\mu})}>0$ for $k=0,\dots,d-1$. The first part for $k<d$ is a well-known fact for orthogonal polynomials, see Szegő [24, Theorem 3.3.1]. We conclude the second part by Lemma 9.3.
@@
    This gives the claim for $k=d$. Let $k\in\{2,\dots,d\}$ and assume as induction hypothesis that
    $x_{k+1-i,k}\leqslant\lambda_{i}^{2},\quad i=1,\dots,k.$
@@
    $x_{(k-1)+1-i,k-1}<x_{k+1-i,k},\quad i=1,\dots,k-1,$
    such that the claim for $k-1$ follows from the induction hypothesis. ∎
The following result is originally due to Nemirovskii [20, (3.14)], who introduced the function $\varphi_{k}^{2}$, named $r_{k}$ therein. The proof of the first inequality can also be found in Hanke [10, (3.8)] and Engl, Hanke and Neubauer [8, (7.7)]. For the sake of completeness, we state it in our notation below. The second inequality is new, but follows easily from 9.6(b) and the definition of $\varphi_{k}$.
@@ 9.3 Proofs for Section 4
  1. 1.
    Note that $r_{k+1}-r_{k}\in\operatorname{Pol}_{k+1,0}$. Lemma 9.4 implies $\langle{\mathsf{r_{k}}},{\mathsf{r_{k+1}}}\rangle_{Y}=R_{k+1}^{2}$. We conclude the claim by
    $R_{t}^{2}=(1-\alpha)^{2}R_{k}^{2}+2(1-\alpha)\alpha\langle{\mathsf{r_{k}}},{\mathsf{r_{k+1}}}\rangle_{Y}+\alpha^{2}R_{k+1}^{2}.$
  2. 2.
[…]
  3. 3.
@@
    Let $k=0$. By (9.4) and $\alpha_{1}<\alpha_{2}$, we have
    $x_{1,t_{2}}=\alpha_{2}^{-1}x_{1,1}<\alpha_{1}^{-1}x_{1,1}=x_{1,t_{1}}.$
    This gives the claim for $k=0$. Now, let $k=1,\dots,d-1$. The definitions of $r_{t_{1}}$ and $r_{t_{2}}$ and (9.2) yield
@@
    since $r_{k+1}(x)<r_{k}(x)$ for all $x\in(0,x_{1,k+1}]$ by 9.6(e), $r_{k}(x)\geqslant 0$ for $x\in[0,x_{1,k}]$ and $r_{k+1}(x)<0$ for $x\in(x_{1,k+1},x_{1,k}]$. This implies $x_{1,t_{2}}<x_{1,t_{1}}$. Let $x>x_{k+1,k+1}$. For $k$ even we obtain $r_{k}(x)>0$ and $r_{k+1}(x)<0$ such that $r_{t_{1}}(x)>r_{t_{2}}(x)$. Since $r_{t_{j}}(x_{k+1,k+1})>0$, $j=1,2$, this yields $x_{k+1,t_{2}}<x_{k+1,t_{1}}$. In the same way, for $k$ odd we deduce $r_{t_{1}}(x)<r_{t_{2}}(x)$ and $r_{t_{j}}(x_{k+1,k+1})<0$, $j=1,2$, implying $x_{k+1,t_{2}}<x_{k+1,t_{1}}$.
@@
    The right-continuity in $t_{0}=0$ is immediate, since we have for $t=\alpha\in(0,1]$ by (9.4)
    $r_{t,<}(x)=\Big{(}1-\alpha\frac{x}{x_{1,1}}\Big{)}{\bf 1}(x<\alpha^{-1}x_{1,1}),\quad x\in[0,\lVert AA^{\top}\rVert_{\mathrm{spec}}].$
[…]
    $\sup_{x\in[0,\lVert AA^{\top}\rVert_{\mathrm{spec}}]}\lvert r_{t_{0},<}(x)-r_{t,<}(x)\rvert$ $\leqslant\sup_{x\in[0,x_{1,t}]}\lvert r_{t_{0}}(x)-r_{t}(x)\rvert$
    $=(\alpha_{0}-\alpha)\sup_{x\in[0,x_{1,t}]}\lvert r_{k+1}(x)-r_{k}(x)\rvert$
    $\leqslant(t_{0}-t)\sup_{x\in[0,\lVert AA^{\top}\rVert_{\mathrm{spec}}]}\lvert r_{k+1}(x)-r_{k}(x)\rvert,$
    where the latter supremum is finite. Letting $t\uparrow t_{0}$ shows the left-continuity for all $t_{0}\in(0,d]$, and right-continuity for $t_{0}\in(0,d)$ can be proven in the same way. This gives the first claim.
@@
and note the order $0\leqslant\tilde{x}_{1,t}<x_{1,k+1}<x_{1,t}$ due to $x_{i,t}>x_{i,k+1}$ by 4.5(b). With some $q_{k},\tilde{q}_{k},\dot{q}_{k}\in\operatorname{Pol}_{k,0}$, $\tilde{p}_{k-1}\in\operatorname{Pol}_{k-1,1}$, we have
@@
$=\frac{\tilde{x}_{1,t}}{x_{1,t}}+\dot{q}_{k}(x)-\frac{x}{x_{1,t}}\bigg{(}x^{k}\prod_{i=2}^{k+1}\Big{(}-\frac{1}{x_{i,t}}\Big{)}+\tilde{p}_{k-1}(x)\bigg{)}$
@@
$\leqslant\Big{(}\frac{2\mu+1}{e\lvert r_{t}^{\prime}(0)\rvert}\Big{)}^{2\mu+1}\lVert g\rVert_{\mu+1/2}^{2}$
@@
Derivation of (4.2).
[…]
$\alpha_{\lambda}\coloneqq{}$ $\operatorname*{argmin}_{\alpha\in{\mathbb{R}}}\Big{(}\lVert\mathsf{r_{k}}+\alpha(\mathsf{r_{k+1}-r_{k}})\rVert_{Y}^{2}+\lambda a_{k+1}(r_{k+1})^{2}\alpha^{2}\Big{)}$
[…]
where we used that $\langle{\mathsf{r_{k+1}-r_{k}}},{\mathsf{r_{k+1}}}\rangle_{Y}=0$ by Lemma 9.4. The numerator is just $R_{k}^{2}-R_{k+1}^{2}$, which is strictly positive for $k+1\leqslant d$ so that $\alpha_{\lambda}\in(0,1]$. ∎